OpenStudy (anonymous):

find the exact value for cos^-1(cos(-5pi/3))

OpenStudy (anonymous):

Do the stuff inside the parenthesis first. cos(-5π/3) = -1/2 Then work your way outside.

OpenStudy (anonymous):

x = cos^-1(-1/2) cos(x) = -1/2 x = 2π/3 and 4π/3

OpenStudy (anonymous):

\[\cos^-1*\cos(5\pi/3)=1.8388\]

OpenStudy (anonymous):

the answer key says the answer is pi/3 but I don't know how to do it

OpenStudy (anonymous):

okay lets do it this way

OpenStudy (anonymous):

Hey I miscalculated there, cos(-5π/3) = 1/2

OpenStudy (anonymous):

cosx = 1/2 => x = pi/3 or -pi/3

OpenStudy (anonymous):

usually for a function \(f^{-1}\circ f(x)=x\) but not in this case, because \(-\frac{5\pi}{3}\) is not in the range of \(\cos^{-1}(x)\) the range of \(\cos^{-1}(x)\) is \([0,\pi]\)

OpenStudy (anonymous):

so what you are looking for is a number (angle) on the upper half of the unit circle where cosine is the same as it is at \(-\frac{5\pi}{3}\)

OpenStudy (anonymous):



OpenStudy (anonymous):

that is clearly coterminal with \(\frac{\pi}{3}\) and so that is your answer

OpenStudy (anonymous):

btw the answer is NOT \(\frac{\pi}{3}\) or \(-\frac{\pi}{3}\) arccosine is a well defined function

OpenStudy (anonymous):

thanks bro